OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -17, -13, -9, -5, ...

OpenStudy (sagewilson):

What pattern do you see happening?

OpenStudy (anonymous):

increasing

OpenStudy (sagewilson):

Yes, increasing by how much?

OpenStudy (anonymous):

4

OpenStudy (sagewilson):

Yeah, I'll give you the list until you get to positive numbers. -17, -13, -9, -5, -1, 3, +4, +4, +4 <-- answer

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

whats next

OpenStudy (sagewilson):

Just keep adding 4, start with 3 + 4 then do the sum of that plus 4 and repeat that 2 more times.

OpenStudy (anonymous):

sigh

OpenStudy (anonymous):

you kknow that says NTH not NINTH

OpenStudy (anonymous):

right

OpenStudy (sagewilson):

.....No I did not realize that. Give me a minute.

OpenStudy (anonymous):

this is precalc

OpenStudy (sagewilson):

Okay I haven't done PreCalc so I had to do some research. An equation I'm getting is "Nth term = A1+ (An - 1)D" Since -17 is the first term and 4 is the difference it would be Nth = -17 (An - 1)4. Still I don't know where the An - 1 comes in.

OpenStudy (sagewilson):

Alright An is the term you're looking for but they aren't asking for a term just Nth.

OpenStudy (sagewilson):

Okay so An would be the numbers they've given you, so it would be Nth = -17 + (4 - 1)4. -17 + (3)4. -17 + 12 = -5. So I think your Nth term would be -5.